OpenStudy (anonymous):

PLEASE HELP!!! URGENT!! WILL GIVE MEDAL! A charge of 6.7 × 10^-15 coulombs is located at a point where its potential energy is 5.6 × 10^-12 joules. What is the electric potential at that point? A. 2.3 × 10^2 volts B. 4.7 × 10^2 volts C. 8.4 × 10^2 volts D. 9.2 × 10^2 volts

OpenStudy (anonymous):

Look up the formula that relates electric potential and potential energy.

OpenStudy (anonymous):

U = 5.6 × 10^-12 J q = 6.7 × 10^-15 C U = qV V = U/q = 836 V So, the correct answer is C

OpenStudy (anonymous):

Oh my god!! THANK YOU!

OpenStudy (anonymous):

@SimonRiley You're welcome